OpenStudy (anonymous):

Please Help. How do I figure this out? A roofer requires 10 h to shingle a roof. After the roofer and an apprentice work on a roof for 5 h, the roofer moves on to another job. The apprentice requires 15 more hours to finish the job. How long would it take the apprentice, working alone, to do the job?

OpenStudy (p0sitr0n):

1 roof = 10 hours 15x + (x+y)5 = 1 x =1/10 y = 1/40 40 hours

OpenStudy (anonymous):

thanks